a) If Maria’s utility function is U = A Xayb what is his marginal utility functions for X and Y? Find the MRS between x and y.
b) What is his marginal rate of substitution between X =3 and Y=4? If A=1, a = 0.4 and b = 0.5.
1
Expert's answer
2020-07-15T10:03:14-0400

a) "MRSxy=\frac{MUx}{MUy}"


"MUx=\frac{\delta U}{\delta X}=AaX^{a-1}Y^b"


"MUy=\frac{\delta U}{\delta Y}=AbX^aY^{b-1}"


"MRSxy=\frac{MUx}{MUy}=\frac{AaX^{a-1}Y^b}{AbX^aY^{b-1}}=\frac{aX^{a-1}Y^b}{bX^aY^{b-1}}"

b)"MRSxy=\frac{MUx}{MUy}=\frac{AaX^{a-1}Y^b}{AbX^aY^{b-1}}=\frac{aX^{a-1}Y^b}{bX^aY^{b-1}}=\frac{0.4\times3^{0.4-1}4^{0.5}}{0.5\times3^{0.4}\times4^{0.5-1}}=1.067"